OpenStudy (anonymous):

Simplify 8x+6y/2xy please show steps and explain

OpenStudy (saifoo.khan):

\[\frac{2(4x+3y)}{2xy}\] \[\frac{\cancel 2(4x+3y)}{\cancel 2xy}\] \[\frac{4x+3y}{xy}\]

OpenStudy (anonymous):